Back to the NIKA2 main page

Daily Reports

Special session after the new sweep (Oct 16-17; MC) inside a heterodyne Pool (Manager: Mónica Rodríguez). Times referred to LT.

Oct 18th, Monday

05:10; Ángel: Opacity (tau[225 GHz] from radiometer) about 0.4, relatively stable. NIKA2 pointing model set to Summer. Bad kids, total kids after new sweep and start acquisition (window open): 22, 1055 (arr1); 5, 573 (arr2); 33, 1056 (arr3). Some problems at the beginning with imbfits creation in ~/mnt/data/NIKAIMBFITS/current/. Fixed: OK. Setting to project nikas-21. Pointing on W3OH. A message of corrupted file/synchro problems from piic-monitor2pf appeared (due to kids geometry changes, I guess). IDL pipeline configured to run66 instead of run67. Adopting pointing 0.,0. and focus -0.8. Beammap on Uranus (scan 34). Slight opacity spike in the middle. Ephemeris fluxes from ASTRO: 19.1 Jy/b (150 GHz) and 44.9 Jy/b (260 GHz). Then a skydip (scan 35). Changing focus to -1.0. A second beammap on Uranus was tried at 07:50 (scan 38) with a mean opacity ~0.45 +/- 0.10. End of time agreed. Additional notes in TAPAS.

Oct 19th, Tuesday

22:30-10:00; Ioannis, Miguel, Ángel: EMIR receiver halted during this night because of an apparent problem with the IF switch box. Taking advantage of this situation we switched to NIKA2. Opacity about 0.2 from midnight with some isolated instabilities. Switching to old sweep (run66.ini) at the beginning. Acquisition started. Setting to project nikas-21. From a calib1scan (scan 162) on Uranus we got p2p fluctuations of +/-4 Jy/b (2 mm) and +/-15 Jy/b (1 mm); mean beam sizes of 18.4 and 12.8 arcsec (2 and 1 mm). After pointing/focus on Uranus we installed the new sweep (run67.ini) to get a beam map on that source (20211019s176) at El. ~45 deg. Two additional beam maps were obtained by following this procedure, corresponding to the scans 20211020s7 at El. ~66 deg., option h, and s33 at El. ~54 deg. For the latter one, the opacity was about 0.25 +/- 0.10. A skydip with the new sweep was also obtained (20211020s8). We started science observations of the Pool on behalf of one of the low-rated projects: 083-21. We got the provided scan series on targets J0311+0507, J0839+0113, and J1018+1117.

DailyReportsNika2Pool181021 (last edited 2021-10-20 08:02:24 by NikaBolometer)